# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

### v1

@@ Appendix C Comparison of Large Language Models for CTN Caption Generation
Appendix D Impact of Automatic Evaluation on CTN Caption Generation
@@
Similarly, in Figure 15, the CTN caption generated with automatic evaluation correctly identifies the cause of the protests in Brazil as the increase in bus ticket prices and links it to the effect of a woman being interviewed about her participation in the protests. The caption generated without automatic evaluation, while mentioning the protest and its cause, fails to capture the specific effect of the woman’s interview.
Figure 16 demonstrates how the CTN caption generated with automatic evaluation accurately describes the cause of a terror attack in Texas and its effect on news anchors discussing freedom of speech issues. The caption generated without automatic evaluation is completely unrelated, mentioning a band’s performance and the audience’s reaction.
@@
Appendix E Application of CTN Caption Generation for Labeling Unlabeled Videos
Figure 18: Application of our CTN caption generation approach for labeling unlabeled videos. Given an unlabeled video, we extract frames and generate image captions using a state-of-the-art image captioning model (GIT [51]). These captions are then used as input to our LLM-based CTN caption generation pipeline, which produces a CTN caption for the entire video. The generated CTN caption captures the cause-effect relationships and temporal dynamics in the video, enabling effective labeling of unlabeled video content.
@@ Appendix E Application of CTN Caption Generation for Labeling Unlabeled Videos
Figure 18 illustrates the pipeline for labeling an unlabeled video using our CTN caption generation approach. First, we extract frames from the unlabeled video. In this example, we extract five frames which are equally spaced in the video. Next, we generate image captions for each of the extracted frames using a state-of-the-art image captioning model. For this demonstration, we employ the GIT [51], which has shown impressive performance in generating accurate captions for individual images. The GIT model generates captions such as "a tractor is driving", "there are four trucks", "a tractor is upside down", "a game is being played", and "two people are celebrating" for the five frames. These image captions serve as the input to our LLM-based CTN caption generation pipeline (see Figure 2).
@@
To ensure the quality and relevance of the generated CTN caption, we employ the same evaluation framework described in Section 3.1. The caption is compared against the video content using the EMScore [41] metric, and only captions that meet a specified threshold are retained. This application demonstrates the versatility and effectiveness of our CTN caption generation approach in labeling unlabeled videos. By leveraging the power of state-of-the-art image captioning models and our few-shot based prompt, we can generate high-quality CTN captions that accurately capture the causal-temporal narrative in video content, even in the absence of human-annotated captions. This approach has the potential to significantly streamline the process of labeling large-scale video datasets and enable more effective video understanding and retrieval tasks.
Appendix F Additional Qualitative Results for CEN Architecture

### v2

@@ A.3 Comparison of Large Language Models for CTN Caption Generation
A.4 Impact of Automatic Evaluation on CTN Caption Generation
@@
Similarly, in Figure 15, the CTN caption generated with automatic evaluation correctly identifies the cause of the protests in Brazil as the increase in bus ticket prices and links it to the effect of a woman being interviewed about her participation in the protests. The caption generated without automatic evaluation, while mentioning the protest and its cause, fails to capture the specific effect of the woman’s interview. Figure 16 demonstrates how the CTN caption generated with automatic evaluation accurately describes the cause of a terror attack in Texas and its effect on news anchors discussing freedom of speech issues. The caption generated without automatic evaluation is completely unrelated, mentioning a band’s performance and the audience’s reaction.
@@
A.5 Human Evaluation Study Details
To validate the quality of our CTN captions, we conduct a rigorous human evaluation study. This section details the methodology, results, and analysis of this evaluation.
A.5.1 Methodology
We employ the following methodology for our human evaluation:
  • Sample size: 100 validation instances (50 each from MSR-VTT and MSVD datasets)
  • Evaluators: Five independent domain experts
  • Evaluation criteria:
    1. 1.
      Causal Accuracy: Degree of correctly identifying and describing cause-effect relationships
    2. 2.
      Temporal Coherence: Accuracy in representing the sequence of events
    3. 3.
      Relevance: How well the caption reflects the overall content and context
  • Rating scale: 5-point Likert scale (0-5, with 5 being the highest quality)
A.5.2 Results
Table 3 presents the results of our human evaluation study.
Table 3: Human Evaluation Results for CTN Captions
Criterion Mean Score Std Dev % Perfect Scores
Causal Accuracy 4.8 0.42 82%
Temporal Coherence 4.7 0.46 78%
Relevance 4.9 0.31 91%
Overall Quality 4.8 0.40 84%
Key findings from the evaluation:
  • Overall mean quality score: 4.8/5.0 ($\sigma=0.40$)
  • 93% of captions received scores of 4 or higher across all dimensions
  • 84% of captions achieved perfect scores
A.5.3 Inter-rater Reliability
To ensure the consistency of ratings across evaluators, we calculate the Intraclass Correlation Coefficient (ICC) for absolute agreement among raters:
  • ICC: 0.87 (95% CI: 0.83-0.91)
This high ICC value indicates strong inter-rater reliability, supporting the robustness of our evaluation process.
A.5.4 Analysis
The results of our human evaluation study strongly validate the quality and consistency of our CTN captions. With a mean quality score of 4.8/5.0 and 93% of captions receiving high scores across all dimensions, we can confidently assert that our CTN captions effectively capture causal-temporal narratives in videos. The high inter-rater reliability further strengthens the credibility of these results.
These findings demonstrate that our automatic generation and evaluation process, as described in Section 3.1, produces high-quality CTN captions that accurately represent the causal and temporal dynamics in video content. This human evaluation complements our automatic evaluation metrics, providing a comprehensive validation of our CTN caption benchmark.
A.6 Encoder Comparison for CEN
To explore the versatility of our CEN architecture, we conduct experiments using BLIP Li et al. (2022b) encoders in place of CLIP encoders. This comparison allows us to assess the adaptability of our method to different visual encoding architectures. Table 4 presents the results of this comparison on both MSR-VTT and MSVD datasets.
Table 4: Comparison of CLIP and BLIP Encoders for our CEN
Dataset Encoder ROUGE-L CIDEr SPICE
MSR-VTT CLIP (Ours) 27.90 49.87 15.76
BLIP 25.34 43.01 14.89
MSVD CLIP (Ours) 31.46 63.51 19.25
BLIP 28.79 57.23 18.06
While the performance with CLIP encoders is higher, the results with BLIP encoders are still competitive. This demonstrates the flexibility of our CEN architecture and its potential to work effectively with various visual encoding methods.
A.7 Comparison with NExT-QA
To highlight the differences between our CTN benchmark and the NExT-QA dataset, we provide a detailed comparison table and illustrative examples:
Table 5 presents a comprehensive comparison between NExT-QA and our NarrativeBridge approach.
Table 5: Comparison of NExT-QA and NarrativeBridge
Aspect NExT-QA NarrativeBridge (Ours)
Task Question-Answering Video Captioning
Output Short answers Coherent narratives
Temporal Specific points Multiple events
Causal Implicit Explicit
Narrative Not present Explicit
Focus Info retrieval Video description
Figure 18: Examples of question-answer pairs from the NExT-QA dataset.
This comparison highlights the fundamental differences in approach and output between NExT-QA and our NarrativeBridge method.
Figure 18 provides examples from the NExT-QA dataset, illustrating the nature of its question-answer pairs. As seen in these examples, NExT-QA focuses on specific question-answer pairs that often target single events or actions. In contrast, our CTN captions provide a more comprehensive narrative that captures the causal and temporal relationships across the entire video sequence.
A.8 Application of CTN Caption Generation for Labeling Unlabeled Videos
@@
Figure 19: Application of our CTN caption generation approach for labeling unlabeled videos. Given an unlabeled video, we extract frames and generate image captions using a state-of-the-art image captioning model (GIT Wang et al. (2022)). These captions are then used as input to our LLM-based CTN caption generation pipeline, which produces a CTN caption for the entire video. The generated CTN caption captures the cause-effect relationships and temporal dynamics in the video, enabling effective labeling of unlabeled video content.
Figure 19 illustrates the pipeline for labeling an unlabeled video using our CTN caption generation approach. First, we extract frames from the unlabeled video. In this example, we extract five frames which are equally spaced in the video. Next, we generate image captions for each of the extracted frames using a state-of-the-art image captioning model. For this demonstration, we employ the GIT Wang et al. (2022), which has shown impressive performance in generating accurate captions for individual images. The GIT model generates captions such as "a tractor is driving", "there are four trucks", "a tractor is upside down", "a game is being played", and "two people are celebrating" for the five frames. These image captions serve as the input to our LLM-based CTN caption generation pipeline (see Figure 2).
@@
To ensure the quality and relevance of the generated CTN caption, we employ the same evaluation framework described in Section 3.1. The caption is compared against the video content using the EMScore Shi et al. (2022) metric, and only captions that meet a specified threshold are retained. This application demonstrates the versatility and effectiveness of our CTN caption generation approach in labeling unlabeled videos. By leveraging the power of state-of-the-art image captioning models and our few-shot based prompt, we can generate high-quality CTN captions that accurately capture the causal-temporal narrative in video content, even in the absence of human-annotated captions. This approach has the potential to significantly streamline the process of labeling large-scale video datasets and enable more effective video understanding and retrieval tasks.
A.9 Additional Qualitative Results for CEN Architecture